# Topology-guided Hypergraph Transformer Network: Unveiling Structural Insights for Improved Representation

### v2

@@ Abstract.
Hypergraphs, with their capacity to depict high-order relationships, have emerged as a significant extension of traditional graphs. Although graph neural networks (GNNs) perform remarkably well in graph representation learning, their extension to hypergraphs encounters challenges due to their intricate structures. Furthermore, current hypergraph transformers, a special variant of GNN, utilize semantic feature-based self-attention, ignoring topological attributes of nodes and hyperedges. To address these challenges, we propose a Topology-guided Hypergraph Transformer Network (THTN). In this model, we first formulate a hypergraph from a graph while retaining its structural essence to learn higher-order relations within the graph. Then, we design a simple yet effective structural and spatial encoding module to incorporate the topological and spatial information of the nodes into their representation. Further, we present a structure-aware self-attention mechanism that discovers the important nodes and hyperedges from both semantic and structural viewpoints. By leveraging these two modules, THTN crafts an improved node representation, capturing both local and global topological expressions. Extensive experiments conducted on node classification tasks demonstrate that the performance of the proposed model consistently exceeds that of the existing approaches.
@@ 1. Introduction
While Graph Neural Networks (GNNs) show great performance in graph representation learning, the complexity of hypergraph structure poses a significant hurdle in adapting these models for efficiently learning hypergraph representation. In response, researchers have introduced various hypergraph representation learning techniques, including hypergraph neural networks (Feng et al., 2019), hypergraph convolution networks (Bai et al., 2021), and hypergraph attention networks (Chen et al., 2020a). These methods aim to learn node and hyperedge embeddings by considering both the topology and attributes of hypergraphs, offering improved solutions for understanding complex relationships and patterns within such data. Furthermore, drawing on the success of transformers in graph data analysis, the hypergraph transformer model has been developed to extend this approach to hypergraphs (Li et al., 2023; Xia et al., 2022; Zhang et al., 2022a; Ding et al., 2023). However, current hypergraph transformer models mainly focus on attribute-based semantic features in the self-attention layer, resulting in the loss of the structural attributes of nodes and hyperedges.
@@
To address these limitations, we propose a novel Topology-guided Hypergraph Transformer Network (THTN) model. THTN generates the representation of nodes via a new structure-aware self-attention mechanism that effectively identifies the importance of nodes and hyperedges from both semantic and structural perspectives. As the initial step, we create hypergraphs from standard graphs by considering nodes’ higher-order relations that preserve the structure information in the graph. Then, we define our THTN model on the constructed hypergraph. In the input layer of THTN  we introduce a learnable structure encoding scheme comprising local structure encoding, centrality encoding, and uniqueness encoding to capture important local and global structural information. Furthermore, we utilize a learnable hypergraph Laplacian eigenvector in the input layer as node positional information. This Laplacian eigenvector-based position encoder enables the encoding of distance-aware spatial information within the hypergraph.
@@
Moreover, while a hyperedge consists of multiple nodes, each node possesses varying degrees of structural and semantic importance for that hyperedge. Similarly, a node may belong to multiple hyperedges, but not all hyperedges hold equal structural and semantic importance for that node. To capture individual importance from both structural and semantic perspectives, we introduce a structure-aware self-attention mechanism that consists of two layers. i) Local Structure-Aware Node-to-Hyperedge Attention layer aggregates node representations into a hyperedge representation via emphasizing structurally and semantically significant nodes. To define the structural importance of a node for a hyperedge, we use two measures, node-local clustering coefficient, and node coreness. These measures calculate how densely connected the node’s local neighborhood is. ii) Global Structure-Aware Hyperedge-to-Node Attention layer aggregates hyperedge representations into a node representation via highlighting structurally and semantically important hyperedges. To define the structural importance of a hyperedge for a node, we introduce two new measures, the hyperedge density score and the hyperedge clustering coefficient. Contributions of our work are summarized as follows:
@@
  • Structure-Aware Self-Attention: We introduce a structure-aware self-attention mechanism in addition to the semantic feature-based vanilla self-attention utilized in existing hypergraph transformers. Our attention mechanism in THTN incorporates both attribute-based semantic features and structural information that enables the identification of crucial nodes for a hyperedge and important hyperedges for a node from both points of view. We further introduce four different measures to discover the structurally important nodes and hyperedges.
@@ 2.1.1. Transformer
The transformer is a neural network architecture that exploits a self-attention mechanism to capture local and global dependencies in the input data. The main components of the model are a multi-head self-attention (MHA) module and a position-wise feed-forward network (FFN). The MHA module computes attention weights by projecting the input into a query ($Q$), key ($K$), and value ($V$) vectors. If $Z\in\mathbb{R}^{n\times d}$ is a $d$ dimensional input feature where $Z=[z_{1},z_{2},....,z_{n}]^{T}$, the MHA first projects it to Q, K, and V using three learnable weight matrices $W^{Q}\in\mathbb{R}^{d\times d_{Q}}$, $W^{K}\in\mathbb{R}^{d\times d_{K}}$, and $W^{V}\in\mathbb{R}^{d\times d_{V}}$ as
@@
Then in each head $h\in\{1,2,3,...,H\}$, the self-attention mechanism is applied to the corresponding ($Q_{h},K_{h},V_{h}$) as
@@ 2.1.2. Hypergraph
Figure 1. Topology-guided Hypergraph Transformer Network, THTN, consists of a) Structural and Spatial Encoding that enrich initial node representation via learnable structural and spatial node features and b) Structure-Aware Self-Attention that enables the integration of structural importance of nodes for a hyperedge and hyperedges for a node into regular attribute-based semantic attention to derive the ultimate node representations.
@@ 2.2. Related Work
Vaswani et al. introduce the transformer model in (Vaswani et al., 2017), emphasizing self-attention mechanisms’ ability to capture long-range dependencies in sequential data. This inspired the development of graph transformers, extending self-attention to handle graph-structured data. Dwivedi and Bresson pioneered in this domain by proposing a graph transformer (Dwivedi and Bresson, 2020) that incorporates graph connectivity as an inductive bias, ensuring nodes attend only to their neighbors. This model utilizes the graph Laplacian eigenvectors for node positional encoding. Gophormer (Zhao et al., 2021) takes a different approach, extracting ego-graphs as transformer inputs instead of processing the entire graph. Challenges in assimilating structural data into the transformer model are addressed through the proposal of a proximity-enhanced attention mechanism (Zhao et al., 2021) and by incorporating centrality, spatial, and edge encoding techniques (Ying et al., 2021).
While transformers for standard graphs have seen considerable progress, a gap remains in developing transformers specifically for hypergraphs. Recent research in hypergraph transformers (Li et al., 2023; Liu et al., 2023; Xia et al., 2022; Zhang et al., 2022b; Ding et al., 2023) shows promise but tends to focus predominantly on the attribute-based semantic features of nodes and hyperedges within the self-attention module, while often overlooking crucial structural information. Notably, models in (Li et al., 2023; Liu et al., 2023) are limited to meta-path-guided heterogeneous hypergraphs, further narrowing their scope. This study addresses this gap by introducing a topology-guided transformer model for hypergraphs that consists of a structural and spatial encoding module and a structure-aware self-attention mechanism that considers both the semantic and structural importance of nodes and hyperedges. By incorporating both aspects, the proposed model provides a more comprehensive representation and understanding of the underlying data.
@@ 3. Methodology
In this section, we outline components of our THTN model. As the first step of THTN, we present the structure and spatial encoding module as the initial node feature enrichment process. Next, we propose a structure-aware self-attention module that integrates both the local and global structural information of nodes and hyperedges into the regular semantic feature-based self-attention network as an inductive bias.
@@ 3.1.4. Position Encoding
3.2. Structure-Aware Self-Attention
In this section, we describe our structure-aware self-attention mechanism capturing high-order relationships between nodes and hyperedges. This model incorporates novel two-level attention, encompassing both node-to-hyperedge and hyperedge-to-node information propagation via considering their semantic and structural importance for each other, facilitating the integration of local and global contexts into node representation learning.
The first level, node-to-hyperedge, aggregates node information to create hyperedge representations via local structure-aware self-attention. Conversely, the second level, hyperedge-to-node, gathers hyperedge representations to generate the final node representation via global structure-aware self-attention. As shown in Equation 2, the self-attention network in traditional transformer architecture calculates attention scores between different nodes and hyperedges by evaluating the attribute-based semantic similarity whereas the vital structural information is ignored. To address this, at each attention level, we introduce different structural inductive biases that allow the model to capture both semantic and vital structural information simultaneously.
@@ 3.2. Structure-Aware Self-Attention
Hyperedge is degree-free and consists of an arbitrary number of nodes. However, the contribution of nodes and their importance in hyperedge construction could be different. To identify the important nodes for a given hyperedge, we leverage a self-attention mechanism that aggregates node representations and assigns higher weights to crucial ones. In addition to a semantic perspective of regular self-attention, we inherit the local structure of subgraphs as hyperedges to extract the structural significance of nodes within hyperedges. Since we represent each subgraph as a hyperedge, the structurally significant nodes for a subgraph will also be important for that corresponding hyperedge. To determine these nodes, we calculate the structural importance of nodes for that subgraph (hyperedge). Then, we combine nodes’ structural importance with the attention map as a structure inductive bias. With the modified attention mechanism, the $l$-th layer representation $q_{j}^{l}$ of a hyperedge $e_{j}$ is defined as
@@
where $\alpha$ is a nonlinear activation function, $W_{1}$ is a trainable weight matrix, and $\Gamma_{ji}$ is the attention coefficient of node $v_{i}$ in the hyperedge $e_{j}$. The attention coefficient is defined as
@@
In hypergraphs, a node can be part of several hyperedges, yet not all hyperedges hold equal significance for a node. This necessitates a self-attention mechanism to emphasize the key hyperedges associated with a specific node. Moreover, employing vanilla self-attention as defined in Equation 2 will only capture the importance of hyperedges from a semantic point of view, ignoring structural importance. To address this issue, we dive into the structural properties of hyperedges, and the corresponding subgraphs and calculate the structural importance of the subgraphs (hyperedges) for each node.
@@
where $\alpha$ is a nonlinear activation function, $W_{4}$ is a trainable weight matrix, $E_{i}$ is the set of hyperedges connected to node $v_{i}$, and $\Lambda_{ij}$ is the attention coefficient of hyperedge $e_{j}$ on node $v_{i}$. The attention coefficient is defined as
@@
Like the transformer paper (Vaswani et al., 2017), the output of each self-attention layer is fed into a position-wise FFN layer. Moreover, residual connections and layer normalization are used to stabilize and normalize the output. THTN generates the node representations by utilizing these two levels of attention. Finally, the output of THTN is linearly projected with a shared trainable weight matrix $W_{S}$ to generate an $S$ dimensional output for each node as $O=pW_{S}$, where $S$ is the number of classes, $p$ is the output of the THTN. We train our entire model using a cross-entropy loss function.
@@ 4.1. Experimental Setup
We evaluate THTN on five distinct datasets including citation network and social network datasets (Fey and Lenssen, 2019). Detailed statistics for these datasets are provided in Table 1. The community detection algorithm follows (Chen et al., 2010) with default settings. We also present constructed hypergraph statistics of these datasets in Table 1. The number of global nodes for the Cora and the Citeseer, PubMed, DBLP, and LastFMAsia datasets is set to 3, 1, 4, 5, and 4, respectively. Our experiments follow a standard split of 50% for training, 25% for validation, and 25% for testing.
To assess performance, we benchmark THTN against sixteen models using their default settings. These models are categorized based on their architectural design: models intended for standard graphs are referred to as graph-based models, while those designed for hypergraphs are referred to as hypergraph-based models. The baseline hypergraph-based models utilize hypergraphs constructed according to the methodologies described in their respective original studies. The graph-based models include Graph Transformer (GT) full and sparse versions (Dwivedi and Bresson, 2020), Graphormer (Ying et al., 2021), ANS-GT (Zhang et al., 2022c), NAGphormer (Chen et al., 2022), GPRGNN (Chien et al., 2021b), GCNII (Chen et al., 2020b). The hypergraph-based models include HGNN (Feng et al., 2019), HCHA (Bai et al., 2021), HyperGCN (Yadati et al., 2019), DHGNN (Jiang et al., 2019), HNHN (Dong et al., 2020), UniGCNII (Huang and Yang, 2021), AllSetTransformer (Chien et al., 2021a), SheafHyperGCN and SheafHyperGNN (Duta et al., 2024), HSL (Cai et al., 2022), and DHKH (Kang et al., 2022).
@@
Method Model Cora Citeseer PubMed DBLP LastFMAsia
Graph-based GT-full 63.10$\pm$0.52 59.05$\pm$0.78 77.02$\pm$1.01 78.85$\pm$0.25 79.23$\pm$0.42
GT-sparse 72.04$\pm$0.30 66.12$\pm$0.85 80.11$\pm$0.78 82.15$\pm$0.15 81.21$\pm$1.23
Graphormer 70.76$\pm$1.52 67.52$\pm$0.56 82.17$\pm$0.56 80.01$\pm$0.65 83.23$\pm$1.45
ANS-GT 86.70$\pm$0.82 73.52$\pm$0.55 87.92$\pm$0.88 82.52$\pm$1.65 86.88$\pm$0.63
NAGphormer 86.12$\pm$0.72 71.55$\pm$1.15 87.51$\pm$0.38 81.89$\pm$1.27 86.15$\pm$1.22
GPRGNN 86.17$\pm$0.68 76.11$\pm$0.22 86.96$\pm$0.68 81.78$\pm$0.46 85.92$\pm$0.85
GCNII 86.71 $\pm$ 0.82 75.86 $\pm$ 0.25 88.07 $\pm$ 0.65 81.58 $\pm$ 0.78 85.96 $\pm$ 0.60
Hypergraph-based HGNN 79.39$\pm$1.36 72.45$\pm$1.16 86.44 $\pm$0.44 78.25 $\pm$0.25 80.48 $\pm$0.64
HCHA 79.14 $\pm$ 1.02 72.42 $\pm$ 1.42 86.41 $\pm$ 0.36 79.52 $\pm$0.80 82.44 $\pm$1.79
HyperGCN 78.45 $\pm$ 1.26 71.28 $\pm$ 0.82 82.84 $\pm$ 8.67 82.78 $\pm$1.23 80.20 $\pm$1.56
DHGNN 79.52$\pm$1.19 73.59$\pm$0.50 80.50 $\pm$ 1.35 80.37 $\pm$0.62 80.22 $\pm$0.80
HNHN 76.36 $\pm$ 1.92 72.64 $\pm$1.57 86.90 $\pm$ 0.30 80.72 $\pm$1.26 84.17 $\pm$1.15
UniGCNII 78.81$\pm$1.05 73.05$\pm$2.21 88.25 $\pm$ 0.40 82.17 $\pm$0.36 84.49 $\pm$1.54
AllSetTransformer 78.59$\pm$1.47 73.08$\pm$1.20 88.72 $\pm$ 0.37 83.15 $\pm$0.56 86.21 $\pm$1.07
SheafHyperGCN 80.06 $\pm$ 1.12 73.27 $\pm$ 0.50 87.09 $\pm$ 0.71 82.37 $\pm$ 0.12 86.88 $\pm$ 0.32
SheafHyperGNN 81.30 $\pm$ 1.70 74.71 $\pm$ 1.23 87.68 $\pm$ 0.60 83.11 $\pm$ 0.23 87.14 $\pm$ 0.55
HSL 79.88$\pm$0.73 73.79$\pm$1.62 - - -
DHKH 82.60 $\pm$0.55 71.68 $\pm$1.20 77.50 $\pm$0.80 81.48 $\pm$0.82 83.25 $\pm$1.56
THTN 88.48$\pm$1.10 78.09$\pm$0.6 89.01$\pm$0.3 84.38$\pm$1.2 88.58$\pm$1.35
@@
Table 3. Impact of local and global structural knowledge on the model performance (accuracy) for only local structures (LS) (first row) and only global structures (GS) (second row).
THTN Cora Citeseer PubMed DBLP LastFMAsia
LS 83.12 73.89 85.32 81.13 83.86
GS 79.68 70.56 82.88 79.32 80.36
@@
4.2. Result
We evaluate the performance of our model by conducting experiments on five distinct datasets and comparing the results with thirteen state-of-the-art baselines. The baselines are considered if their experimental results or codes are available. The outcomes, presented in Table 2, unambiguously demonstrate the superiority of our model across all datasets. Specifically, our model excels on the Cora dataset, achieving an impressive accuracy of $88.48\%$. This significantly surpasses the accuracy of the best-performing graph-based baseline model, GCNII, at $86.71\%$ and exceeds the top-reported accuracy of the hypergraph-based baseline, DHKH, which stands at $82.60\%$. In the case of the Citeseer dataset, our model attains an accuracy of $78.09\%$, outperforming the graph-based leading baseline GPRGNN with an accuracy of $76.11\%$, and the hypergraph-based top-performing baseline, SheafHyperGNN, at $74.71\%$. The trend continues with the PubMed, DBLP, and LastFMAsia datasets, where our model substantially outperforms the baselines. The results underscore our model’s substantial enhancements in classifying the datasets, setting a new standard compared to existing state-of-the-art methods.
Table 4. Comparative impact of different design schemes on the model performance (accuracy) for Cora and PubMed datasets.

### v4

@@ Abstract.
Hypergraphs, with their capacity to depict high-order relationships, have emerged as a significant extension of traditional graphs. Although graph neural networks (GNNs) perform remarkably well in graph representation learning, their extension to hypergraphs encounters challenges due to their intricate structures. Furthermore, current hypergraph transformers, a special variant of GNN, utilize semantic feature-based attention, ignoring topological attributes of nodes and hyperedges. To address these challenges, we propose a Topology-guided Hypergraph Transformer Network (THTN). In this model, we first formulate a hypergraph from a graph while retaining its structural essence to learn higher-order relations within the graph. Then, we design a simple yet effective structural and spatial encoding module to incorporate the topological and spatial information of the nodes into their representation. Further, we present a structure-aware attention mechanism that discovers the important nodes and hyperedges from both semantic and structural viewpoints. By leveraging these two modules, THTN crafts an improved node representation, capturing both local and global topological expressions. Extensive experiments conducted on node classification tasks demonstrate that the performance of the proposed model consistently exceeds that of the existing approaches.
@@ 1. Introduction
While Graph Neural Networks (GNNs) show great performance in graph representation learning, the complexity of hypergraph structure poses a significant hurdle in adapting these models for efficiently learning hypergraph representation. In response, researchers have introduced various hypergraph representation learning techniques, including hypergraph neural networks (Feng et al., 2019), hypergraph convolution networks (Bai et al., 2021), and hypergraph attention networks (Chen et al., 2020a). These methods aim to learn node and hyperedge embeddings by considering both the topology and attributes of hypergraphs, offering improved solutions for understanding complex relationships and patterns within such data. Furthermore, drawing on the success of transformers in graph data analysis, the hypergraph transformer model has been developed to extend this approach to hypergraphs (Li et al., 2023; Xia et al., 2022; Zhang et al., 2022a; Ding et al., 2023). However, current hypergraph transformer models mainly focus on attribute-based semantic features in the attention layer, resulting in the loss of the structural attributes of nodes and hyperedges.
@@
To address these limitations, we propose a novel Topology-guided Hypergraph Transformer Network (THTN) model. THTN generates the representation of nodes via a new structure-aware attention mechanism that effectively identifies the importance of nodes and hyperedges from both semantic and structural perspectives. As the initial step, we create hypergraphs from standard graphs by considering nodes’ higher-order relations that preserve the structure information in the graph. Then, we define our THTN model on the constructed hypergraph. In the input layer of THTN  we introduce a learnable structure encoding scheme comprising local structure encoding, centrality encoding, and uniqueness encoding to capture important local and global structural information. Furthermore, we utilize a learnable hypergraph Laplacian eigenvector in the input layer as node positional information. This Laplacian eigenvector-based position encoder enables the encoding of distance-aware spatial information within the hypergraph.
@@
Moreover, while a hyperedge consists of multiple nodes, each node possesses varying degrees of structural and semantic importance for that hyperedge. Similarly, a node may belong to multiple hyperedges, but not all hyperedges hold equal structural and semantic importance for that node. To capture individual importance from both structural and semantic perspectives, we introduce a structure-aware attention mechanism that consists of two layers. i) Local Structure-Aware Node-to-Hyperedge Attention layer aggregates node representations into a hyperedge representation via emphasizing structurally and semantically significant nodes. To define the structural importance of a node for a hyperedge, we use two measures, node-local clustering coefficient, and node coreness. These measures calculate how densely connected the node’s local neighborhood is. ii) Global Structure-Aware Hyperedge-to-Node Attention layer aggregates hyperedge representations into a node representation via highlighting structurally and semantically important hyperedges. To define the structural importance of a hyperedge for a node, we introduce two new measures, the hyperedge density score and the hyperedge clustering coefficient. Contributions of our work are summarized as follows:
@@
  • Structure-Aware Attention: We introduce a structure-aware attention mechanism in addition to the semantic feature-based vanilla attention utilized in existing hypergraph transformers. Our attention mechanism in THTN incorporates both attribute-based semantic features and structural information that enables the identification of crucial nodes for a hyperedge and important hyperedges for a node from both points of view. We further introduce four different measures to discover the structurally important nodes and hyperedges.
@@ 2.1.1. Transformer
The transformer is a neural network architecture that exploits an attention mechanism to capture local and global dependencies in the input data. The main components of the model are a multi-head self-attention (MHA) module and a position-wise feed-forward network (FFN). The MHA module computes attention weights by projecting the input into a query ($Q$), key ($K$), and value ($V$) vectors. If $Z\in\mathbb{R}^{n\times d}$ is a $d$ dimensional input feature where $Z=[z_{1},z_{2},....,z_{n}]^{T}$, the MHA first projects it to Q, K, and V using three learnable weight matrices $W^{Q}\in\mathbb{R}^{d\times d_{Q}}$, $W^{K}\in\mathbb{R}^{d\times d_{K}}$, and $W^{V}\in\mathbb{R}^{d\times d_{V}}$ as
@@
Then in each head $h\in\{1,2,3,...,H\}$, the attention mechanism is applied to the corresponding ($Q_{h},K_{h},V_{h}$) as
@@ 2.1.2. Hypergraph
Figure 1. Topology-guided Hypergraph Transformer Network, THTN, consists of a) Structural and Spatial Encoding that enrich initial node representation via learnable structural and spatial node features and b) Structure-Aware Attention that enables the integration of structural importance of nodes for a hyperedge and hyperedges for a node into regular attribute-based semantic attention to derive the ultimate node representations.
@@ 2.2. Related Work
Vaswani et al. introduce the transformer model in (Vaswani et al., 2017), emphasizing attention mechanisms’ ability to capture long-range dependencies in sequential data. This inspired the development of graph transformers, extending attention to handle graph-structured data. Dwivedi and Bresson pioneered in this domain by proposing a graph transformer (Dwivedi and Bresson, 2020) that incorporates graph connectivity as an inductive bias, ensuring nodes attend only to their neighbors. This model utilizes the graph Laplacian eigenvectors for node positional encoding. Gophormer (Zhao et al., 2021) takes a different approach, extracting ego-graphs as transformer inputs instead of processing the entire graph. Challenges in assimilating structural data into the transformer model are addressed through the proposal of a proximity-enhanced attention mechanism (Zhao et al., 2021) and by incorporating centrality, spatial, and edge encoding techniques (Ying et al., 2021).
While transformers for standard graphs have seen considerable progress, a gap remains in developing transformers specifically for hypergraphs. Recent research in hypergraph transformers (Li et al., 2023; Liu et al., 2023; Xia et al., 2022; Zhang et al., 2022b; Ding et al., 2023) shows promise but tends to focus predominantly on the attribute-based semantic features of nodes and hyperedges within the attention module, while often overlooking crucial structural information. Notably, models in (Li et al., 2023; Liu et al., 2023) are limited to meta-path-guided heterogeneous hypergraphs, further narrowing their scope. This study addresses this gap by introducing a topology-guided transformer model for hypergraphs that consists of a structural and spatial encoding module and a structure-aware attention mechanism that considers both the semantic and structural importance of nodes and hyperedges. By incorporating both aspects, the proposed model provides a more comprehensive representation and understanding of the underlying data.
@@ 3. Methodology
In this section, we outline components of our THTN model. As the first step of THTN, we present the structure and spatial encoding module as the initial node feature enrichment process. Next, we propose a structure-aware attention module that integrates both the local and global structural information of nodes and hyperedges into the regular semantic feature-based attention network as an inductive bias.
@@ 3.1.4. Position Encoding
3.2. Structure-Aware Attention
In this section, we describe our structure-aware attention mechanism capturing high-order relationships between nodes and hyperedges. This model incorporates novel two-level attention, encompassing both node-to-hyperedge and hyperedge-to-node information propagation via considering their semantic and structural importance for each other, facilitating the integration of local and global contexts into node representation learning.
The first level, node-to-hyperedge, aggregates node information to create hyperedge representations via local structure-aware attention. Conversely, the second level, hyperedge-to-node, gathers hyperedge representations to generate the final node representation via global structure-aware attention. As shown in Equation 2, the attention network in traditional transformer architecture calculates attention scores between different nodes and hyperedges by evaluating the attribute-based semantic similarity whereas the vital structural information is ignored. To address this, at each attention level, we introduce different structural inductive biases that allow the model to capture both semantic and vital structural information simultaneously.
@@ 3.2. Structure-Aware Attention
Hyperedge is degree-free and consists of an arbitrary number of nodes. However, the contribution of nodes and their importance in hyperedge construction could be different. To identify the important nodes for a given hyperedge, we leverage an attention mechanism that aggregates node representations and assigns higher weights to crucial ones. In addition to a semantic perspective of regular attention, we inherit the local structure of subgraphs as hyperedges to extract the structural significance of nodes within hyperedges. Since we represent each subgraph as a hyperedge, the structurally significant nodes for a subgraph will also be important for that corresponding hyperedge. To determine these nodes, we calculate the structural importance of nodes for that subgraph (hyperedge). Then, we combine nodes’ structural importance with the attention map as a structure inductive bias. With the modified attention mechanism, the $l$-th layer representation $q_{j}^{l}$ of a hyperedge $e_{j}$ is defined as
@@
where $\alpha$ is a nonlinear activation function, $W_{1}$ is a trainable weight matrix, and $\Gamma_{ji}$ is the attention coefficient of node $v_{i}$ in the hyperedge $e_{j}$.
@@
In hypergraphs, a node can be part of several hyperedges, yet not all hyperedges hold equal significance for a node. This necessitates an attention mechanism to emphasize the key hyperedges associated with a specific node. Moreover, employing vanilla attention as defined in Equation 2 will only capture the importance of hyperedges from a semantic point of view, ignoring structural importance. To address this issue, we dive into the structural properties of hyperedges, and the corresponding subgraphs and calculate the structural importance of the subgraphs (hyperedges) for each node.
@@
where $\alpha$ is a nonlinear activation function, $W_{4}$ is a trainable weight matrix, $E_{i}$ is the set of hyperedges connected to node $v_{i}$, and $\Lambda_{ij}$ is the attention coefficient of hyperedge $e_{j}$ on node $v_{i}$.
@@
Like the transformer paper (Vaswani et al., 2017), the output of each attention layer is fed into a position-wise FFN layer. Moreover, residual connections and layer normalization are used to stabilize and normalize the output. THTN generates the node representations by utilizing these two levels of attention. Finally, the output of THTN is linearly projected with a shared trainable weight matrix $W_{S}$ to generate an $S$ dimensional output for each node as $O=pW_{S}$, where $S$ is the number of classes, $p$ is the output of the THTN. We train our entire model using a cross-entropy loss function.
@@ 4.1. Experimental Setup
We evaluate THTN on six distinct datasets including citation network and social network datasets (Fey and Lenssen, 2019). Detailed statistics for these datasets are provided in Table 1. The community detection algorithm follows (Chen et al., 2010) with default settings. We also present constructed hypergraph statistics of these datasets in Table 1. The number of global nodes for the Cora and the Citeseer, PubMed, DBLP, LastFMAsia, and Wiki datasets is set to 3, 1, 4, 5, 4, and 4 respectively. Our experiments follow a standard split of 50% for training, 25% for validation, and 25% for testing.
To assess performance, we benchmark THTN against eighteen models using their default settings. These models are categorized based on their architectural design: models intended for standard graphs are referred to as graph-based models, while those designed for hypergraphs are referred to as hypergraph-based models. The baseline hypergraph-based models utilize hypergraphs constructed according to the methodologies described in their respective original studies. The graph-based models include Graph Transformer (GT) full and sparse versions (Dwivedi and Bresson, 2020), Graphormer (Ying et al., 2021), ANS-GT (Zhang et al., 2022c), NAGphormer (Chen et al., 2022), GPRGNN (Chien et al., 2021b), GCNII (Chen et al., 2020b). The hypergraph-based models include HGNN (Feng et al., 2019), HCHA (Bai et al., 2021), HyperGCN (Yadati et al., 2019), DHGNN (Jiang et al., 2019), HNHN (Dong et al., 2020), UniGCNII (Huang and Yang, 2021), AllSetTransformer (Chien et al., 2021a), SheafHyperGCN and SheafHyperGNN (Duta et al., 2024), HSL (Cai et al., 2022), and DHKH (Kang et al., 2022).
@@
Method Model Cora Citeseer PubMed DBLP LFMA Wiki
Graph-based GT-full 63.10$\pm$0.52 59.05$\pm$0.78 77.02$\pm$1.01 78.85$\pm$0.25 79.23$\pm$0.42 68.13$\pm$0.88
GT-sparse 72.04$\pm$0.30 66.12$\pm$0.85 80.11$\pm$0.78 82.15$\pm$0.15 81.21$\pm$1.23 72.44$\pm$0.68
Graphormer 70.76$\pm$1.52 67.52$\pm$0.56 82.17$\pm$0.56 80.01$\pm$0.65 83.23$\pm$1.45 73.11$\pm$1.05
ANS-GT 86.70$\pm$0.82 73.52$\pm$0.55 87.92$\pm$0.88 82.52$\pm$1.65 86.88$\pm$0.63 76.24$\pm$1.08
NAGphormer 86.12$\pm$0.72 71.55$\pm$1.15 87.51$\pm$0.38 81.89$\pm$1.27 86.15$\pm$1.22 76.55$\pm$0.75
GPRGNN 86.17$\pm$0.68 76.11$\pm$0.22 86.96$\pm$0.68 81.78$\pm$0.46 85.92$\pm$0.85 74.64$\pm$0.32
GCNII 86.71$\pm$0.82 75.86$\pm$0.25 88.07$\pm$0.65 81.58$\pm$0.78 85.96$\pm$0.60 74.66$\pm$0.58
Hypergraph-based HGNN 79.39$\pm$1.36 72.45$\pm$1.16 86.44$\pm$0.44 78.25$\pm$0.25 80.48$\pm$0.64 73.82$\pm$0.92
HCHA 79.14$\pm$1.02 72.42$\pm$1.42 86.41$\pm$0.36 79.52 $\pm$0.80 82.44 $\pm$1.79 74.13$\pm$1.32
HyperGCN 78.45$\pm$1.26 71.28$\pm$0.82 82.84$\pm$8.67 82.78 $\pm$1.23 80.20$\pm$1.56 74.22$\pm$1.34
DHGNN 79.52$\pm$1.19 73.59$\pm$0.50 80.50$\pm$1.35 80.37 $\pm$0.62 80.22$\pm$0.80 75.52$\pm$0.68
HNHN 76.36$\pm$1.92 72.64$\pm$1.57 86.90$\pm$0.30 80.72$\pm$1.26 84.17$\pm$1.15 75.14$\pm$0.92
UniGCNII 78.81$\pm$1.05 73.05$\pm$2.21 88.25$\pm$0.40 82.17 $\pm$0.36 84.49$\pm$1.54 77.18$\pm$1.22
AllSetTransformer 78.59$\pm$1.47 73.08$\pm$1.20 88.72$\pm$0.37 83.15 $\pm$0.56 86.21$\pm$1.07 78.49$\pm$0.82
SheafHyperGCN 80.06$\pm$1.12 73.27$\pm$0.50 87.09$\pm$0.71 82.37$\pm$ 0.12 86.88$\pm$0.32 77.42$\pm$0.72
SheafHyperGNN 81.30$\pm$1.70 74.71$\pm$1.23 87.68$\pm$0.60 83.11 $\pm$0.23 87.14$\pm$0.55 78.17$\pm$0.92
HSL 79.88$\pm$0.73 73.79$\pm$1.62 - - - -
DHKH 82.60$\pm$0.55 71.68$\pm$1.20 77.50$\pm$0.80 81.48$\pm$0.82 83.25$\pm$1.56 75.18$\pm$1.23
THTN 88.48$\pm$1.10 78.09$\pm$0.6 89.01$\pm$0.3 84.38$\pm$1.2 88.58$\pm$1.35 79.15$\pm$0.55
@@
Table 3. Impact of local and global structural information on the model performance (accuracy).
THTN Cora Citeseer PubMed DBLP LFMA Wiki
LS 83.12 73.89 85.32 81.13 83.86 75.23
GS 79.68 70.56 82.88 79.32 80.36 72.22
@@
4.2. Comparison with Baseline Methods
We evaluate the performance of our model by conducting experiments on six distinct datasets and comparing the results with eighteen state-of-the-art baselines. The baselines are considered if their experimental results or codes are available. The outcomes, presented in Table 2, unambiguously demonstrate the superiority of our model across all datasets. Specifically, our model excels on the Cora dataset, achieving an impressive accuracy of $88.48\%$. This significantly surpasses the accuracy of the best-performing graph-based baseline model, GCNII, at $86.71\%$ and exceeds the top-reported accuracy of the hypergraph-based baseline, DHKH, which stands at $82.60\%$. In the case of the Citeseer dataset, our model attains an accuracy of $78.09\%$, outperforming the graph-based leading baseline GPRGNN with an accuracy of $76.11\%$, and the hypergraph-based top-performing baseline, SheafHyperGNN, at $74.71\%$. The trend continues with the PubMed, DBLP, LastFMAsia (LFMA), and Wiki datasets, where our model substantially outperforms the baselines. The results underscore our model’s substantial enhancements in classifying the datasets, setting a new standard compared to existing state-of-the-art methods.
Figure 3. The performance (accuracy) of THTN with different numbers of global nodes ($n_{g}$).
Table 4. Impact of different design schemes on the model performance (accuracy) for Cora and PubMed datasets.